# WallFacer: Guiding Transformer Model Training Out of the Long-Context Dark Forest with N-body Problem

### v2

@@ 4.1. Overview
Built upon the inspiration from n-body simulation, the WallFacer training system is comprised of five main components. At the core of the system is WallFacer Attention, which leverages multiple ring-style P2P (peer-to-peer) communication strategies to enhance the efficiency of distributed attention computation. The Dataloader is designed to organize tokens within each sub-sequence according to their mask (causal or full) and distribute them across different GPUs. The Communication Configuration generator plays a critical role in initially assigning Keys and Values to their corresponding ranks. Meanwhile, the Communication Topology Scheduler outlines the placement of parallelism across various computing nodes. Finally, the WallFacer Runtime incorporates additional supporting techniques for the training process, such as gradient checkpointing. We will now explore the details of these components in depth.
4.2. WallFacer Attention
Figure 4. An example of a Transformer block with WallFacer Attention on one team of four devices out of 64 GPUs.
@@ 4.2. WallFacer Attention
As discussed in the previous section, a major limitation of Ring Attention is the extensive amount of peer-to-peer (P2P) communication required, which becomes problematic in environments with weak connections between computing nodes. Drawing inspiration from the multi-dimensional parallelism used in n-body simulations, we enhance the ring sequence parallelism by introducing an additional dimension. This is achieved by duplicating the Queries, Keys, and Values, thus dividing the communication tasks and distributing them within each team. This segmentation markedly improves computational efficiency and scalability.
@@
We will now delve into the specifics of the WallFacer Parallelism and provide a theoretical analysis demonstrating its advantages over Ring Attention.
4.2.1. Training Process of WallFacer Parallelism
@@
In the WallFacer system, GPUs are grouped into Teams to coordinate computation and communication tasks more effectively, unlike the isolated operations in traditional settings. The Queries, Keys, and Values are gathered within the team, so each member of the team holds the same activation of the whole team before attention. WallFacer introduces an additional parameter, C, which determines the replication factor of the input and, consequently, the number of GPUs within each team. The range of C is from 1 to $\sqrt{P}$. When C equals one, the algorithm falls back to Ring Attention. When C equals $\sqrt{P}$, the algorithm becomes a completely collective-communication-based one. When $1<C<\sqrt{P}$, it becomes a structure with multiple rings looping concurrently. There are three main distinctions between the operational processes of WallFacer and Ring Attention: QKV_matmul & all-gather, attention iteration, and reduce-scatter. Throughout the attention process, asynchronous communication is employed alongside the early launch of communication kernels to maximize the overlap of computation and communication tasks. The complete forward process of a WallFacer Transformer block is outlined in Algorithm 1.
@@
0:  Input sequence x, Linear Function query, key, and value, attention parallelism size c, global rank r, global size gs, team process group pg
@@
2:  compute the initial rank to send $\textbf{r}_{send}=$$get\_init\_send(\textbf{r})$
3:  compute the initial rank to receive from $\textbf{r}_{recv}=$$get\_init\_recv(\textbf{r})$
@@
6:  initialize attention score O, extra statistics lse to zero.
7:  for $1\leq i\leq world\_size/c^{2}$ do
8:     wait for $\textbf{req}_{send}$ and $\textbf{req}_{recv}$
9:     $\textbf{k}_{current}=\textbf{k}_{next}$, $\textbf{v}_{current}=\textbf{v}_{next}$
10:     launch $\textbf{req}_{send}$ to send $\textbf{k}_{current}$ and $\textbf{v}_{current}$ to $\textbf{r}_{next}$, launch $\textbf{req}_{recv}$ to receive $\textbf{k}_{next}$ and $\textbf{v}_{next}$ from $\textbf{r}_{last}$
[…]
12:  end for
[…]
14:  return $\textbf{O}_{final}$
@@
Once this phase is complete, each GPU within the team possesses the same Q, K, and Vs, each of a length of $\frac{CN}{P}$. To distribute the communication and computation tasks among the team members, we divide the original workload based on four specific ranks assigned to each GPU by the Communication Configuration Generator. These ranks determine each GPU’s partners and position within the P2P ring, details of which will be discussed in Section 4.3.
@@
Backward Propagation. The major distinction between backward and forward propagation, as outlined in section 3, is the inability to calculate Queries independently during the backward phase. Unlike forward propagation, the backward phase requires the complete set of Keys and Values to calculate the gradient for Queries, and vice versa. To manage this, we have structured the gradient calculation into two loops: the Key & Value outer loop and the Query inner loop.
In the outer loop, gradients for Keys and Values are tracked and maintained fixed on the corresponding GPUs within the sub-rings; these gradients do not transfer between GPUs. The inner loop, however, handles the gradients for Queries, which start initialized as zero and are circulated along the sub-rings together with the Queries themselves. During each iteration, the approach mirrors the backward computation method used in Flash Attention, where the updated gradient of the current Query shard is passed to the next GPU in the ring, while the gradients for Keys and Values are retained for subsequent Query shards.
@@
4.2.2. Theoretical Analysis
In this section, we will discuss two major questions: how much communication reduction can WallFacer bring compared with Ring Attention, and does this come with a high extra memory cost? During the analysis, we will employ a case study using the WallFacer system with an attention parallel size of C = 4 on a llama-30B model, which consists of 64 layers. For this model, referred to as model M, the batch size = B is set to 1, the sequence length = N to 65,536, the hidden dimension = H to 6,656, and the number of GPUs = P to 64. Additionally, the computation will utilize bfloat16 precision.
Communication Analysis. Let’s analyze the communication overhead within one forward Transformer block on a single GPU. For Ring Attention, the communication is primarily due to the ring P2P loop. Each iteration’s communication overhead can be calculated as follows:
(4) $\frac{2BNH}{PW}+L$
and as the total number of iterations done is $P$, the total communication overhead will be:
(5) $\frac{2BNH}{W}+PL$
@@ 4.2.2. Theoretical Analysis
For WallFacer, the communication overhead comes from collective and P2P both. The collective overhead for all-gather and reduce-scatter is:
(6) $\frac{4BNH(C-1)}{PW}$
@@
(7) $\frac{P}{C^{2}}(\frac{2CBNH}{PW}+L)=\frac{2BNH}{CW}+\frac{P}{C^{2}}L$
@@
Memory Analysis. In this section, we estimate the theoretical peak memory requirements necessary to store the model weights, activations, and optimizer states. Our implementation utilizes the Adam Optimizer (Kingma and Ba, 2017), bfloat16 precision, and Zero-2 optimization (Rajbhandari et al., 2020). Since the WallFacer architecture does not alter the model weights or the optimizer, we can assume that the memory costs associated with the model and the optimizer remain constant across both methods. We name the memory cost for the model and optimizer as $M_{m+o}$. As for the activation, we refer to the size of one single activation of a sub-sequence on one GPU as
(8) $A=\frac{B\times N\times H}{P}$
As we use the checkpointing scheme from (Li et al., 2024), a model of $Y$ layers needs to save $Y+1$ activations as checkpoints. Now we calculate the approximate peak memory after Q, K, and V are already calculated and before the attention computation at the last layer of the whole model. For Ring Attention and WallFacer, the peak memories are:
(9) $PM_{Ring}=M_{m+o}+(L+1)A+3A=M_{m+o}+(L+4)A$
(10) $PM_{Wall}=M_{m+o}+(L+1)A+3CA=M_{m+o}+(L+3C+1)A$
@@
4.3. Communication Configuration Generator
The Communication Topology Scheduler, as introduced in the previous section, plays a crucial role in the initial setup of Queries, Keys, and Values on GPUs and determining each GPU’s position within the sub-rings.
Figure 5. An example of ring initialization process of 8GPUs and 4 sub-rings in the Communication Configuration Generator
@@ 4.3. Communication Configuration Generator
Furthermore, Team 0 and Team 1 form what is termed a team group, where ring P2P communication occurs exclusively within this group, similarly for Team 2 and Team 3. A team group is defined as a collection of teams that participate in the same sub-rings. The number of teams within a team group can be calculated as $Team\ size=\frac{P}{C^{2}}$. We give the details of the determination of initial sending targets in the case of placing teams within same computing nodes in Algorithm 2, and the receiving source can also be calculated similarly.
@@
4.4. Communication Topology Scheduler
@@
When using Ring Attention, all GPUs form a single ring, which considerably limits the flexibility of adjusting the placement scheme. In contrast, WallFacer expands the tuning space with its unique multi-ring structure in two significant ways. First, the ring configuration in WallFacer is determined by the parallelism dimension $C\in[1,\sqrt{P}]$. Smaller $C$ results in less collective communication, a higher total communication volume, and more inter-node communication, while higher C leads to more communication volume during all-gather and reduce-scatter, which may exceed the overhead of the QKV matrix multiplication that we use for overlapping, but also highly reduce the overall communication volume.
@@
(11) $\begin{split}Config=\arg\max_{C,placement}(Profile(C\in[1,\sqrt{P}],\\ placement\in[P2P\_intra,Collect\_intra])\end{split}$
@@
Figure 6. A comparison of using naive and zigzag dataloader for 8 GPUs with attention parallel dimension of 2. The corresponding initialization can be found in Figure 5 with the same configuration. The improvement of efficiency from load-balancing increases with the number of GPUs.
4.5. Sequence Parallelism Dataloader
To accommodate both full mask and causal mask configurations for self-attention, we have implemented two distinct data loading schemes for load-balancing. For full masks, sequences are straightforwardly divided into equal sub-sequences and distributed among the devices. However, causal masks present a challenge due to the unbalanced computational load across GPUs; sub-sequences at the beginning of a sequence require significantly more computation than those at the end. To address this imbalance and achieve load equilibrium among GPUs, we modify the ZigZag scheme introduced by (Zhu et al., 2024), illustrated in Figure 6. The figure illustrates the simplest case of zigzag load-balancing. Notably, the effectiveness of this strategy improves as the number of GPUs increases. This improvement correlates with the expanding difference in computation volume between the first and the last token, which escalates as the sequence length extends. This approach ensures that the total workload on each GPU is balanced, eliminating the need for additional communication mechanisms like those employed in DistFlashAttention(Li et al., 2024).
4.6. WallFacer Implementation
WallFacer is written in PyTorch(Paszke et al., 2019) and uses the PyTorch torch.autograd.function and NCCL(NVIDIA, 2020) backend for forward and backward implementation. WallFacer also employs multiple techniques during runtime to improve its overall training efficiency.
@@ 4.6. WallFacer Implementation
Save recomputation with checkpoints. WallFacer adopts the checkpointing strategy introduced by DistFlashAttn(Li et al., 2024), placing checkpoints at the end of the self-attention phase rather than the FFN of each transformer layer. This checkpoint placement effectively obviates the need to recompute the self-attention forward process during the backward pass, avoiding redundant attention computation.
5. Evaluation
Figure 7. Throughput evaluation of Ring Attention and WallFacer on 32 GPUs from three different clusters. We place the performance of WallFacer with both C=2 and C=4 in the figure. The configurations are marked in the titles of the sub-figures. For instance, A100_8(3B, 63K) represents that the experiment is on machines with 8 Nvidia A100 GPUs in each node, the model used has three billion parameters, and the sequence length is 64k.
The computational resources we use in the experiments include a local computing cluster with eight nodes, each equipped with eight Nvidia H100 GPUs with 80GB of high-bandwidth memory, interconnected by NVLink and linked between nodes via InfiniBand with eight Mellanox ConnectX-7 controllers each. Additionally, we utilize Google Cloud Virtual Computing Machines in two setups: one with two machines containing 16 Nvidia A100 GPUs, and another with four machines containing eight GPUs each, both configurations using 40GB of high-bandwidth memory per GPU. All machines are connected within via NVLink and between by Ethernet.
We utilize two different model sizes for the two different GPU types. For H100 with 80GB HBM, we use a GPT model with approximately 7 billion parameters (referred to as 7B in the following text), which has 32 layers, 32 attention heads, and a hidden dimension of 4096. As for the A100 with 40 GB HBM, we use a smaller model with around 3 billion parameters (referred to as 3B), which has 12 layers, 16 attention heads, and a hidden dimension of 4096. During training, both models use bfloat16 precision and a batch size of 1 to accommodate longer input sequences.

### v3

@@ 3.1 Overview
The WallFacer training system is comprised of five main components. At the core of the system is WallFacer Attention, which leverages multiple ring-style P2P (peer-to-peer) communication strategies to enhance the efficiency of distributed attention computation. The Communication Configuration Generator plays a critical role in initially assigning keys and values to their corresponding ranks. Meanwhile, the Communication Topology Scheduler outlines the placement of parallelism across various computing nodes. The Dataloader is designed to organize tokens within each sub-sequence according to their mask (causal or full) and distribute them across different GPUs. Finally, the WallFacer Runtime incorporates additional supporting techniques for the training process, such as gradient checkpointing. We will now explore the details of these components in depth.
3.2 WallFacer Attention
Figure 4: An example of WallFacer Attention on 64 GPUs. Each team member forms a sub-ring with members with the same local rank from other teams in the same ring communication group, reducing the communication volume of each team member by 75%.
@@ 3.2 WallFacer Attention
As discussed in the previous section, a major limitation of Ring Attention is the extensive amount of peer-to-peer (P2P) communication required, which becomes problematic in environments with weak connections between computing nodes. We enhance the ring sequence parallelism by introducing an additional dimension. This is achieved by duplicating the queries, keys, and values, thus dividing the communication tasks and distributing them within each team. This segmentation markedly improves computational efficiency and scalability.
@@
We will now go through the specifics of the WallFacer Parallelism and provide a theoretical analysis demonstrating its advantages over Ring Attention.
3.2.1 Training Process of WallFacer Parallelism
@@
In the WallFacer system, GPUs are grouped into Teams to coordinate computation and communication tasks more efficiently, unlike the isolated operations in traditional settings. The queries, keys, and values are gathered within the team, so each member of the team holds the same activation of the whole team before attention. WallFacer introduces an additional parameter, $C$, which determines the replication factor of the input and, consequently, the number of GPUs within each team. The range of $C$ is from 1 to $\sqrt{P}$. When $C$ equals one, the algorithm falls back to Ring Attention. When $C$ equals $\sqrt{P}$, the algorithm becomes a completely collective-communication-based one with no rings. When $1<C<\sqrt{P}$, it becomes a structure with multiple rings looping concurrently. There are three main distinctions between the operational processes of WallFacer and Ring Attention: QKV_matmul & all-gather, attention iteration, and reduce-scatter. Throughout the attention process, asynchronous communication is employed alongside the early launch of communication kernels to maximize the overlap of computation and communication tasks. The complete forward process of a WallFacer Transformer block is outlined in Algorithm 1.
@@
0:  Input sequence x, Linear Function query, key, and value, attention parallelism size c, global rank r, global size gs, team process group pg, init send/recv target $\textbf{r}_{\mathrm{send}}$ and $\textbf{r}_{\mathrm{recv}}$
1:  compute the gathered $\textbf{q}_{\mathrm{team}},\textbf{k}_{\mathrm{team}},\textbf{v}_{\mathrm{team}}=$AllGather_QKVmatmul(query, key, value, x, pg)
@@
3:  get the ring P2P target $\textbf{r}_{\mathrm{next}}$ and $\textbf{r}_{\mathrm{last}}$ withget_P2P_ranks(r, gs, c)
4:  initialize attention score O, extra statistics lse to zero.
5:  for $1\leq i\leq$world_size$/c^{2}$ do
6:     wait for $\textbf{req}_{\mathrm{send}}$ and $\textbf{req}_{\mathrm{recv}}$
7:     $\textbf{k}_{\mathrm{current}}=\textbf{k}_{\mathrm{next}}$, $\textbf{v}_{\mathrm{current}}=\textbf{v}_{\mathrm{next}}$
8:     launch $\textbf{req}_{\mathrm{send}}$ to send $\textbf{k}_{\mathrm{current}}$ and $\textbf{v}_{\mathrm{current}}$ to $\textbf{r}_{\mathrm{next}}$, launch $\textbf{req}_{\mathrm{recv}}$ to receive $\textbf{k}_{\mathrm{next}}$ and $\textbf{v}_{\mathrm{next}}$ from $\textbf{r}_{\mathrm{last}}$
@@
12:  return $\textbf{O}_{\mathrm{final}}$
@@
Once this phase is complete, each GPU within the team possesses the same Q, K, and V, each of a length of $\frac{CN}{P}$. To distribute the communication and computation tasks among the team members, we divide the original workload based on four specific ranks assigned to each GPU by the Communication Configuration Generator. These ranks determine each GPU’s partners and position within the P2P ring, details of which will be discussed in Section 3.3.
@@
Backward Propagation. The major distinction between backward and forward propagation is the inability to calculate queries independently during the backward phase. Unlike forward propagation, the backward phase requires the complete set of keys and values to calculate the gradient for queries, and vice versa. To manage this, we have structured the gradient calculation into two loops: the key & value outer loop and the query inner loop.
In the outer loop, gradients for keys and values are tracked and maintained fixed on the corresponding GPUs within the sub-rings; these gradients do not transfer between GPUs. The inner loop, however, handles the gradients for queries, which start initialized as zero and are circulated along the sub-rings together with the Queries themselves. During each iteration, the approach mirrors the backward computation method used in flash-attention[dao2022flashattention], where the updated gradient of the current query shard is passed to the next GPU in the ring, while the gradients for keys and Values are retained for subsequent query shards.
@@
3.2.2 Theoretical Analysis
In this section, we will discuss two major questions: how much communication reduction WallFacer can bring compared with Ring Attention, and how much extra memory cost it comes with. During the analysis, we will employ a case study using the WallFacer system with an attention parallel size of $C$ = 4 on a llama-30B model, which consists of 64 layers. For this model, referred to as model M, the batch size = $B$ is set to 1, the sequence length = $N$ to 65, 536, the hidden dimension = $H$ to 6, 656, and the number of GPUs = $P$ to 64. Additionally, the computation will utilize bfloat16 precision.
Communication Analysis. Let’s analyze the communication overhead within one forward Transformer block on a single GPU. For Ring Attention, the communication is primarily due to the ring P2P loop. As the total number of iterations done is $P$, the total communication overhead can be calculated as:
$p(\frac{2BNH}{PW}+L)=\frac{2BNH}{W}+PL$ (2)
@@ 3.2.2 Theoretical Analysis
For WallFacer, the communication overhead comes from both collective and P2P. The collective overhead for all-gather and reduce-scatter is:
$\frac{4BNH(C-1)}{PW}$ (3)
@@
$\frac{P}{C^{2}}(\frac{2CBNH}{PW}+L)=\frac{2BNH}{CW}+\frac{P}{C^{2}}L$ (4)
@@
Memory Analysis. In this section, we estimate the theoretical peak memory requirements necessary to store the model weights, activations, and optimizer states. Our implementation utilizes the Adam Optimizer [kingma2017adam], bfloat16 precision, and Zero-2 optimization [rajbhandari2020zero]. Since the WallFacer architecture does not alter the model weights or the optimizer, we can assume that the memory costs associated with the model and the optimizer remain constant across both methods. We name the memory cost for the model and optimizer as $M_{m+o}$. As for the activation, we refer to the size of one single activation of a sub-sequence on one GPU as
$A=\frac{B\times N\times H}{P}$ (5)
As we use the checkpointing scheme from [li2024distflashattn], a model of $Y$ layers needs to save $Y+1$ activations as checkpoints. Now we calculate the approximate peak memory after Q, K, and V are already calculated and before the attention computation at the last layer of the whole model. For Ring Attention and WallFacer, the peak memories are:
$PM_{Ring}=M_{m+o}+(Y+1)A+3A=M_{m+o}+(Y+4)A$ (6)
$PM_{Wall}=M_{m+o}+(Y+1)A+3CA=M_{m+o}+(Y+3C+1)A$ (7)
@@
3.3 Communication Configuration Generator
The Communication Configuration Generator, as introduced in the previous section, plays a crucial role in the initial setup of queries, keys, and values on GPUs and determining each GPU’s position within the sub-rings.
Figure 5: An example of ring initialization process of 8GPUs and 4 sub-rings in the Communication Configuration Generator
@@ 3.3 Communication Configuration Generator
Furthermore, Team 0 and Team 1 form what is termed a team group, where ring P2P communication occurs exclusively within this group, similarly for Team 2 and Team 3. A team group is defined as a collection of teams that participate in the same sub-rings. The number of teams within a team group can be calculated as $Team\ size=\frac{P}{C^{2}}$. We give the details of the determination of initial sending targets in the case of placing teams within the same computing nodes in Algorithm 2, and the receiving source can also be calculated similarly.
@@
3.4 Communication Topology Scheduler
@@
When using Ring Attention, all GPUs form a single ring, which considerably limits the flexibility of adjusting the placement scheme. In contrast, WallFacer expands the tuning space with its unique multi-ring structure in two significant ways. First, the ring configuration in WallFacer is determined by the parallelism dimension $C\in[1,\sqrt{P}]$. Smaller $C$ results in less collective communication, a higher total communication volume, and more inter-node communication, while higher C leads to more communication volume during all-gather and reduce-scatter, which may exceed the overhead of the QKV matrix multiplication that we use for overlapping, but can highly reduce the overall communication volume.
@@
$\begin{split}Config=\arg\max_{C,placement}(Profile(C\in[1,\sqrt{P}],\\ placement\in[P2P\_intra,Collect\_intra])\end{split}$ (8)
@@
Figure 6: A comparison between naive and zigzag dataloader for 8 GPUs with attention parallel dimension of 2. The corresponding initialization can be found in Figure 5 with the same configuration. The improvement of efficiency from load-balancing increases with the number of GPUs.
3.5 Sequence Parallelism Dataloader
To accommodate both full mask and causal mask configurations for self-attention, we have implemented two distinct data loading schemes for load-balancing. For full masks, sequences are straightforwardly divided into sub-sequences of equal lengths and distributed among the devices. However, causal masks present a challenge due to the unbalanced computational load across GPUs; sub-sequences at the beginning of a sequence require significantly more computation than those at the end. To address this imbalance and achieve load equilibrium among GPUs, we modify the ZigZag scheme introduced by [zilin2024ringflashattention], illustrated in Figure 6. The figure illustrates the simplest case of zigzag load-balancing. Notably, the effectiveness of this strategy improves as the number of GPUs increases. This improvement correlates with the expanding difference in computation volume between the first and the last token, which escalates as the sequence length extends. This approach ensures that the total workload on each GPU is balanced, eliminating the need for additional communication mechanisms like those employed in DistFlashAttention[li2024distflashattn].
Figure 7: Throughput evaluation of Ring Attention and WallFacer on 32 GPUs from three different clusters. We place the performance of WallFacer with both C=2 and C=4 in the figure. The configurations are marked in the titles of the sub-figures. For instance, A100_40GB_8(1B, 512K) represents that the experiment is on machines with 8 Nvidia A100 40GB GPUs in each node, the model used has one billion parameters, and the sequence length is 512k.
3.6 WallFacer Runtime
WallFacer is written in PyTorch[pytorch] and uses the PyTorch torch.autograd.function and NCCL[nccl] backend for forward and backward implementation. WallFacer also employs multiple techniques during runtime to improve its overall training efficiency.
@@ 3.6 WallFacer Runtime
Save recomputation with checkpoints. WallFacer adopts the checkpointing strategy introduced by DistFlashAttn[li2024distflashattn], placing checkpoints at the end of the self-attention phase rather than the FFN of each transformer layer. This checkpoint placement effectively obviates the need to recompute the self-attention forward process during the backward pass, avoiding redundant attention computation.
4 Evaluation
The computational resources we use in the experiments include a local Nvidia H100 cluster with eight nodes and three Nvidia A100 clusters, as listed in table 2. We utilize two model types of total three settings, as listed in table 3. For the DiT(Diffusion Transformer) model, we use similar configurations as those in Stable Diffusion 3[esser2024sd3]. We utilize the backbone Diffusion Transformer only, without other components like the text and image encoders. During training, both models use bfloat16 precision and a batch size of 1 to accommodate longer input sequences.
Table 2: Cluster Configurations. All GPUs are connected with NVLink with computing nodes. #GPU represents local size times the number of nodes.
GPU #GPUs Memory Inter-node
Nvidia H100 8*8 80GB InfiniBand
Nvidia A100 16*2 40GB 100Gbps Ethernet
Nvidia A100 8*4 40GB 100Gbps Ethernet
Nvidia A100 16*2 40GB 100Gbps Ethernet
Table 3: Model Configurations
Model Name #Heads #Layers Hidden Dim
GPT 3B 12 16 4096
GPT 7B 32 32 4096
DiT 1B 24 24 1536